solve the system of equation:
3x=7-y
2y=14-6x

Answer by stanbon(75887)

solve the system of equation:
3x=7-y
2y=14-6x
------
y = -2x+7
y = -3x+7
-----
Substitute for "y" and solve for "x":
-2x+7 = -3x+7
x = 0
---
Solve for "y":
y = -2x + 7
y = -2*0 + 7
y = 7
-------------
Cheers,
Stan H.